You test a population of 10,000 teenagers in Kenya for their genotypes of the sickle cell trait. You find

that 1 in 500 teenagers (200 total) have sickle cell anemia, an autosomal recessive disorder.


(a) What is the predicted frequency of the:

-Recessive allele? Show your work. ?


-Dominant allele? Show your work. ?


(b) What is the predicted frequency of the:


-Homozygous dominant genotype? Show your work.


-Heterozygous/carrier genotype? Show your work.


(c) How many individuals should be:


-Homozygous dominant? Show your work.


-Heterozygous/carriers? Show your work.


1
Expert's answer
2021-03-05T04:14:51-0500

(a)

The frequency of affected children reflects the frequency of homozygous recessive genotype ss:

ss = s2 = 1/500 = 0.002

From here, the frequency of the recessive allele equals:

s = (s2)1/2 = 0.045

Since S + s = 1, the frequency of a dominant allele is:

S = 1 - s = 1 - 0.045 = 0.955

(b)

The predicted frequency of homozygous dominant genotype is:

SS = S2 = 0.9552 = 0.912

The predicted frequency of heterozygous dominant genotype is:

2Ss = 2 × 0.955 × 0.045 = 0.086

(c)

The number of homozygous dominant individuals in the population equals:

N(SS) = 0.912 × 10,000 = 9,120

The number of heterozygous dominant individuals in the population equals:

N(Ss) = 0.086 × 10,000 = 860